Two cars leave an intersection. One car travels north; the other travels east.
When the car traveling north had gone 24 miles, the distance between the cars
was four miles more than three times the distance traveled by the car heading
east. Find the distance between the cars at that time.

Answer by mananth(16946)

Two cars lee an intersection. One car travels north; the other travels east.
When the car traveling north had gone 24 miles, the distance between the cars
was four miles more than three times the distance traveled by the car heading
east. Find the distance between the cars at that time.
Let the car travelling east be at a distance x from the intersection when the other car has traveled 24 miles.
distance between cars = 4+3x
x^2+24^2=(3x+4)^2 Pythagoras theorem
x^2 +24^2= 9x^2+24x+16
8x^+24x-560=0
x^2+3x-70=0
x^2+10x-7x-70=0
x(x+10)-7(x+10)=0
(x+10)(x-7)=0
x=7
x^2+24^2=49+576
625
sqrt 625 = 25
So they are at a distance of 25 miles from each other